=================================================================== Tohban report: 2009-03-16 checked data from: 2009-03-8/00:00 UT to 2009-03-14/24:00 UT =================================================================== Data quality/collection status ============= FGM: A: 3/08-3/14: full data coverage B: 3/08-3/14: full data coverage C: 3/08-3/14: full data coverage D: 3/08-3/14: full data coverage E: 3/08-3/14: full data coverage ESA: A: 3/08-3/14: full data coverage B: 3/08-3/14: data gaps 3/08 00:00 - 3/09 00:30 UT, 3/11 05:45 - 3/12 24:00 UT C: 3/08-3/14: full data coverage D: 3/08-3/14: full data coverage E: 3/08-3/14: full data coverage MOM: A: 3/08-3/14: full data coverage B: 3/08-3/14: data gaps 3/08 00:00 - 3/09 00:30 UT, 3/11 05:45 - 3/12 24:00 UT C: 3/08-3/14: full data coverage D: 3/08-3/14: full data coverage E: 3/08-3/14: full data coverage SST: A: 3/08-3/14: full data coverage B: 3/08-3/14: gap 3/11 16:40 UT -3/13 00:34 UT C: 3/08-3/14: full data coverage D: 3/08-3/14: full data coverage E: 3/08-3/14: full data coverage EFI fields + filter bank: A: 3/08-3/14: full data coverage B: 3/08-3/14: full data coverage C: 3/08-3/14: full data coverage D: 3/08-3/14: full data coverage E: 3/08-3/14: full data coverage SCM fields + filter bank: A: 3/08-3/14: full data coverage B: 3/08-3/14: full data coverage C: 3/08-3/14: full data coverage D: 3/08-3/14: full data coverage E: 3/08-3/14: full data coverage Triggers and burst data: A: 3/08-3/14: look fine; 3/13 AE > 800 nT, multiple bursts sampled B: 3/08-3/14: look fine; C: 3/08-3/14: look fine; 3/13 AE > 800 nT, multiple bursts sampled D: 3/08-3/14: look fine; 3/13 AE > 800 nT, multiple bursts sampled E: 3/08-3/14: look fine; 3/13 AE > 800 nT, multiple bursts sampled ================================= Instrument Ops (Significant events) ------------------------ --------------------------------------------------------------------------------- # 2009/067 - 2009/068 Sunday, March 8 #--------------------------------------------------------------------------------- none #--------------------------------------------------------------------------------- # 2009/068 - 2009/069 Monday, March 9 Week 11 #--------------------------------------------------------------------------------- none #--------------------------------------------------------------------------------- # 2009/069 - 2009/070 Tuesday, March 10 #--------------------------------------------------------------------------------- USNAU_13M THEMIS_B 2.740 2009/069 16:30:00 2009/069 16:50:00 TR2 # SOH check and track [JM] - upload new RTS 17 to change RCS Heater Service 1 duty cycle from 50%/50% to 33%/67% ON/OFF - # load ATS table with pre-shadow ESA HV ramp-down and SST attenuator closing BGS_11M THEMIS_D 9.300 2009/069 18:52:35 2009/069 19:17:35 TR10 # Data dump [JM] - upload IDPU FSW 0x51 - ICCR_JSM81 - enable compression for APID 0x440 packets BGS_11M THEMIS_E 9.200 2009/069 20:28:45 2009/069 20:51:00 TR20 # Data dump [JM] - upload IDPU FSW 0x51 - ICCR_JSM81 - enable compression for APID 0x440 packets BGS_11M THEMIS_A 9.100 2009/069 20:56:00 2009/069 21:18:00 TR40 # Data dump [JM] - disable compression of APID 0x440 packets WHS_18M1 THEMIS_E 2.660 2009/069 21:09:00 2009/069 21:34:00 TR2 # SOH check and track - LRO proficiency test #--------------------------------------------------------------------------------- # 2009/070 - 2009/071 Wednesday, March 11 #--------------------------------------------------------------------------------- BGS_11M THEMIS_B 9.500 2009/070 15:14:45 2009/070 15:26:00 TR9 # Data dump [ME] - transmitter on/off via ATS - no problems after 259 min shadow (05:58 - 10:17) AGO_9M THEMIS_B 9.500 2009/070 15:54:00 2009/070 16:09:00 TR10 # Data dump [ME] - transmitter on/off via ATS HBK_10M THEMIS_B 9.400 2009/070 16:29:00 2009/070 16:40:00 TR9 # Data dump [ME] - transmitter on/off via blind acquisition - post-shadow ESA HV ramp-up and SST attenuator motion enable #--------------------------------------------------------------------------------- # 2009/071 - 2009/072 Thursday, March 12 #--------------------------------------------------------------------------------- none #--------------------------------------------------------------------------------- # 2009/072 - 2009/073 Friday, March 13 #--------------------------------------------------------------------------------- BGS_11M THEMIS_E 9.200 2009/072 20:18:10 2009/072 20:43:00 TR20 # Data dump [GL] - upload new RTS 17 to change RCS Heater Service 1 duty cycle from 50%/50% to 33%/67% ON/OFF - # disable compression of APID 0x440 packets #--------------------------------------------------------------------------------- # 2009/073 - 2009/074 Saturday, March 14 #--------------------------------------------------------------------------------- none